OpenStudy (anonymous):

a driver in a car traveling at a speed of 78 km/h sees a cat 101 m away on the road. How long will it take for the car to accelerate uniformly to a stop in exactly 99m?

OpenStudy (valpey):

\[a_t = c\]\[v_t = \frac{78000m}{3600s} - a_t*t\] \[s_t = 78000*t -(0.5)a_t*t^2\]\[v_0=\frac{78000m}{3600s}; v_f=0;s_0=0;s_f=99m\]

OpenStudy (anonymous):

thanks

OpenStudy (valpey):

And I meant\[s_t=\frac{78000}{3600}t-(0.5)a_t*t^2\]If t is in seconds.

OpenStudy (valpey):

and a is in m/s^2